OpenStudy (anonymous):

I've never seen a5 before!!!!! What is the 7th term of the geometric sequence where a1 = -5 and a5 = -1,280? (3 points) Answer 5,120 -5,120 20,480 -20,480

hartnn (hartnn):

a5 means the 5th term.

OpenStudy (anonymous):

oh:)

hartnn (hartnn):

first find r=common ratio, using the formula, \(\huge a_n=a_1r^{n-1}\) n=5, a1=-5,a5=-1280

OpenStudy (anonymous):

okay:)

OpenStudy (anonymous):

what would i plug in for r into the equation?

hartnn (hartnn):

you need to find 'r' from that equation, thats the only unknown variable.

OpenStudy (anonymous):

n = 7 I thought

hartnn (hartnn):

remember that you have the value of a5. n will be 7 when you'll find the 7th term. now you are finding r, using 5th term, so n= 5 first

OpenStudy (anonymous):

oh alright

OpenStudy (anonymous):

-5 x r ^(5-1)

OpenStudy (anonymous):

-5^4?

hartnn (hartnn):

no.... a5=-1280 so, -1280 = -5 r^4 try to isolate and find r from here.

hartnn (hartnn):

first you can divide by -5 on both sides.

hartnn (hartnn):

can you find r ?

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

-1280/5?

hartnn (hartnn):

r^4 = -1280/ -5 =.... ?

OpenStudy (anonymous):

256

hartnn (hartnn):

yes, r^4 = 256, so r=.... ?

OpenStudy (anonymous):

4294967296

hartnn (hartnn):

:O r^4 = 265 \(\large r = \sqrt[4]{256}=...\) don't do 256^4

OpenStudy (anonymous):

Oh really?? my error one sec

OpenStudy (anonymous):

4 sqrt 256 = 4sqrt of (4^4) = 4

OpenStudy (anonymous):

not quite sure if that's correct

hartnn (hartnn):

yes, r= 4 is correct.

OpenStudy (anonymous):

oh okay good

OpenStudy (anonymous):

ill plug it in

hartnn (hartnn):

now you need 7th term now take n = 7 r= 4 a1 =-5 as always. plug these in \[\huge a_n=a_1r^{n-1}\]

OpenStudy (anonymous):

-5 x 4^ (5-1)

OpenStudy (anonymous):

oh right!!! n= 7 this time

hartnn (hartnn):

yes, because you are finding 7th term

OpenStudy (anonymous):

-5 x 4= -20 i don't know how to plug the whole to the power of "(7-1)"

hartnn (hartnn):

again! don't do -5*4 first!! ^ has greater priority, so do 4^{7-1} = 4^6 first then multiply by 5

OpenStudy (anonymous):

4^(7-1) = 4096

hartnn (hartnn):

yes, 4096*(-5)= ?

OpenStudy (anonymous):

-20480

hartnn (hartnn):

you are \(\huge \color{red}{\checkmark}\)